OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, –2) and is parallel to the line y = –4x + 1? x + 4y = 10 x – 4y = –10 4x – y = –10 4x + y = 10

OpenStudy (dumbcow):

parallel means same slope then plug in the point to solve for y-intercept

OpenStudy (anonymous):

The first step I would do is convert your choices into y-intercept so the equations looks the same. Parallel will mean that they have the same slope (-4x)

OpenStudy (anonymous):

Slope is a number, it is m=-4. Then you can use formula y-y1=m(x-x1), where x1=3, y1=-2